If Cos(x-y),Cos(x),Cos(x+y) are in HP, then 1+cos(y)=?


a)Cosx


b)Cos^2 (x)


c)Sinx


d)Sin^2 (x)


Plz explain..

Grade:12

1 Answers

jitender lakhanpal
12 years ago

Dear chilukari,

2/cosx =1/cos(x-y) + 1/cos(x+y)

2/cosx = 2cosxcosy/(cos2xcos2y-sin2xsin2y)-------1

change sin2y = 1-cos2y  in eqution 1 we get

cos2y - cos2xcosy -sin2x=0           

we can write above equation as cos2y-cosy+sin2xcosy-sin2x=0    writing cos2x=1-sin2x

and subsequently we write 

cosy(cosy-1)+ sin2x(cosy-1)=0

(1-cosy)(cosy+sin2x)=0

so we get   cosy = -sin2x

1+cosy =1-sin2x

1+cosy = cos2 x

so option b) is correct

 

 

 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar: Click here to download the toolbar..  

jitender 

askiitians expert